OpenStudy (anonymous):

Find the distance between the x and y intercepts of 7x-21y=8. Please show your work.

OpenStudy (anonymous):

\[\sqrt{\left(-\frac{8}{21}\right)^2+\left(\frac{8}{7}\right)^2}=\frac{8 \sqrt{10}}{21}=1.205 \]

OpenStudy (anonymous):

7 x - 21 y - 8 = 0 Set x to zero and solve for y. Again with y zero solve for x.

OpenStudy (chriss):

y-intercept -> set x to 0 21y=8 y=8/21 point (0,8/21) x-intercept -> set y to 0 7x=8 x=8/7 point (8/7,0) \[distance=\sqrt{(x _{1}-x _{2})^{2}+(y _{1}-y _{2})^{2}}\] \[=\sqrt{(-\frac{8}{7})^{2}+(\frac{8}{21})^{2}}\] \[\approx1.438\]

OpenStudy (chriss):

1.205 is right... I entered it in my calculator wrong.

OpenStudy (anonymous):

zarkon can i get some help ? on my post

OpenStudy (anonymous):

That can happen. I use Mathematica, a program that, so far, has never made a mistake.

OpenStudy (anonymous):

thanks

OpenStudy (chriss):

yw

OpenStudy (anonymous):

Mathematica rules!